OpenStudy (anonymous):

what is the solution to the equation √(2x+10 )-6=4

OpenStudy (anonymous):

√(2x+10 )-6=4 √(2x+10 )=10 2x+10 = 100 2x=90 x=45

OpenStudy (anonymous):

Is that all?

OpenStudy (anonymous):

Thankyou so much!! :)